# Soap Film Drainage Under Tunable Gravity Using a Centrifugal Thin Film Balance

### v1

@@ III.3 Thickness measurements
In Fig. 4(a), we present typical time evolutions of full thickness profiles $h(r,t)$ in Regime 1, illustrating the coexistence of the two regions until the outer region fully invades the center. In Fig. 4(b), we plot $h_{\rm R}$, $h_{\rm 0}$ and $h_{\rm b}$ as functions of time for the same dataset. Since the thickness profile in the central region remains steady, $h_{\rm 0}$ stays constant up to $t^{\star}$, the time at which the central zone disappears. We observe that both $h_{\rm R}$ and $h_{\rm b}$ decrease over time. Moreover, a proportionality relation is verified between these two quantities: as shown in Fig. 4(c), plotting $h_{\rm b}$ against $h_{\rm R}$ reveals a proportionality with a coefficient of approximately 0.87, which holds even at early times.
In Fig. 5(a), we plot these two quantities for the solution with viscosity $\eta=1$ mPa·s and for all rotation speeds, thus covering both Regimes 1 and 2. The proportionality relation between $h_{\rm b}$ and $h_{\rm R}$ holds at the lowest rotation speeds, with a proportionality coefficient of $0.87\pm 0.01$, obtained from a best fit over all corresponding data. At higher rotation speeds, a deviation from this proportionality appears at early times, when the film is relatively thick. We verified that these behaviors are consistent across all tested viscosities. We define $h_{\rm c}$ as the critical thickness below which the proportionality is recovered, and we find that $h_{\rm c}$ decreases with increasing rotation speed. In Fig. 5(b), we present an occurrence diagram distinguishing experiments that exhibit early-time deviations from proportionality. A comparison with the Regime diagram in Fig. 2(c) reveals a strong correlation: full proportionality is observed primarily in Regime 1, while partial proportionality systematically occurs in Regime 2, as well as in a few high-viscosity cases at the highest rotation speeds within Regime 1.
@@
First, when proportionality is observed, the coefficient of 0.87 is consistent with the value reported by Nierstrasz et al. [nierstrasz1998marginal] for vertical films under standard gravity. This supports the interpretation that TFEs nucleate along the frame edge, detach, and migrate radially inward under effective buoyancy, ultimately settling where the film thickness matches that of the TFEs (see the nucleation along the frame edge in Fig. 1(c), as well as the inward migration indicated by the arrows in panels (b) and (c) of the same figure). This scenario assumes that the TFE migration time is short compared to the overall thinning time, which is always the case in our observations. Our findings validate this mechanism even under centrifugal acceleration up to almost 100 times standard gravity. The gravity-independence of the 0.87 ratio reinforces the scenario proposed by Gros et al. [gros2021marginal], which attributes the selection mechanism to capillary suction and Marangoni stresses, rather than gravitational forces.
@@
which is the equation of an oscillator around the position where $h=h_{\rm TFE}$ as expressed in the third term. The first term corresponds to the inertial force, with a patch mass of $\rho h_{\rm TFE}\pi\ell_{\rm{TFE}}^{2}/4$ and added mass factor $\alpha$. The second term is a damping force with contributions from surface viscosity $\eta_{\rm s}$, bulk viscosity $\eta$, and air friction (with viscosity $\eta_{\rm{air}}$), such that the total two-dimensional viscosity is $\eta^{2D}=2\eta_{\rm s}+\eta h+k_{\rm{air}}\eta_{\rm{air}}\ell_{\rm{TFE}}$, where $k_{\rm{air}}$ is a geometric prefactor [louyer_sliding_2025]. When damping dominates, TFEs directly reach their equilibrium position and $h_{\rm TFE}=h_{\rm b}$. This is consistent with observations in Regime 1 and Regime 2 when $h_{\rm R}<h_{\rm c}$. In contrast, during the early stages of Regime 2 (when $h_{\rm R}>h_{\rm c}$), inertia dominates, and TFEs overshoot their equilibrium and reach the film center.
@@
In experiments, we found that $\ell_{\rm{TFE}}$ can be considered approximately constant (around 2 mm), both at the boundary between the two regions in Regime 1 and at the film center in Regime 2, when $h_{\rm R}$ reaches $h_{\rm c}$ (Fig. 6(b)). This value appears to be independent of $\eta$ and $\omega$, allowing $\ell_{\rm TFE}$ to be treated as constant in Eq. (7). A detailed space–time analysis of $\ell_{\rm TFE}$ is beyond the scope of the present study.
Figure 6: (a) $h_{\rm R}(t=0)$, $h_{\rm c}$ (when a deviation is observed), and $h_{\rm R}(t=t_{\rm f})$ as functions of $\omega$, where $t_{\rm f}$ denotes the final measurement time. Data from all experiments and viscosities are shown with opacity encoding. (b) $\ell_{TFE}$ as a function of $\omega$, showing that this quantity remains approximately constant. Measurements are taken when $h_{\rm R}=h_{\rm c}$, while TFEs are located at the film center. Data include all experiments in which a deviation from proportionality is observed.
@@ III.4 Thinning dynamics
The expression for the linear flux $q$, as given in Eq. (1)—previously validated under standard gravity [lhuissier2012bursting, vigna_flowing_2025, cantat_drainage_2025]—suggests a correlation between the rate of thinning $\textrm{d}h_{\rm R}/\textrm{d}t$ and the film thickness $h_{\rm R}$. Approximating the film volume as $\pi R^{2}h_{\rm R}$ and the liquid loss rate as $2\pi Rq$, we expect, following [lhuissier2012bursting, vigna_flowing_2025, cantat_drainage_2025]:
@@
In Fig. 7(b), we explore the influence of rotation speed by plotting $\lvert\mathrm{d}h_{\rm R}/\mathrm{d}t\rvert$ versus $h_{\rm R}$ at fixed viscosity. Each dataset again follows a power law with an exponent close to $5/2$. Deviations appear when $h_{\rm R}>h_{\rm c}$, as shown by the triangles markers in Fig. 7(c) where the proportionality between $h_{\rm R}$ and $h_{\rm b}$ no longer holds; these data are excluded from the analysis that follows. As expected, higher rotation speeds lead to faster thinning. A power-law fit yields $\lvert\mathrm{d}h_{\rm R}/\mathrm{d}t\rvert\propto\omega^{1.31\pm 0.33}$. Although $\omega$ does not appear explicitly in Eq. (8), it influences the meniscus radius $r_{\rm m}$, which we assume to scale with the effective capillary length, $r_{\rm m}\propto\sqrt{\gamma/(\rho R\omega^{2})}$. Substituting this relation in Eq. (8) gives the thinning law
$\left\lvert\frac{\textrm{d}h_{\rm R}}{\textrm{d}t}\right\rvert\propto\frac{\omega^{3/2}\rho^{3/4}}{\eta\,R^{1/4}}\,h_{R}^{5/2},$ (9)
with $K={\omega^{3/2}\rho^{3/4}}/({\eta\,R^{1/4}})$ the proportionality constant between $\lvert\textrm{d}h_{R}/\textrm{d}t\rvert$ and $h_{R}^{5/2}$. The predicted dependence $K\propto\omega^{3/2}$ agrees well with the empirical exponent $1.31\pm 0.33$. When $\omega^{-3/2}\,\lvert\mathrm{d}h_{\rm R}/\mathrm{d}t\rvert$ is plotted against $h_{\rm R}$ for the data in Fig. 7 (b), all curves collapse onto a single master curve (see inset), confirming the predicted $\omega$ scaling.
Overall, the expression for $K$ is validated over 3 orders of magnitude in Fig. 7(c), where the effects of both viscosity and rotation speed were independently tested. These results confirm that Eq. (1) remains valid even under high-gravity conditions. They are also consistent with the asymptotic behavior $h_{\rm R}\propto t^{-2/3}$ observed at long times (Fig. 4 (b)), in agreement with previous studies under standard gravity [lhuissier2012bursting, miguet2021marginal].
@@ III.5 Discussion
At early stages, in the absence of marginal regeneration, centrifugal acceleration likely induces an initial stretching of the film rather than thinning through internal flow, although existing interfacial elasticity models fail to reproduce the observed profiles. This stretching occurs only in the central region and at low effective gravity, in Regime 1, before erosion by marginal regeneration begins. In Regime 2, TFEs populate the entire film as soon as imaging begins, suggesting that either the initial erosion of the central zone is too fast to be captured, or that the film is sufficiently flat for TFEs to migrate directly toward the center. Interestingly, films remain very stable, particularly in the absence of glycerol. Assuming an average film thickness of a few micrometers, the surface tension difference between the center and the frame edge can be estimated from Eq. (3) as $\Delta\gamma=\rho hR^{2}\omega^{2}/4\sim 7$ mN·m-1, smaller than but comparable to the surface tension of the surfactant solution. This difference may explain why the system approaches the limit of film stability reported in previous studies [de2001young].
@@
Across all Regimes and viscosities, thinning is consistently associated with the nucleation of TFEs in the vicinity of the meniscus. These elements are thinner than the surrounding film, resulting in a net mass loss as they migrate inward. On average, we find a proportionality factor of $0.87\pm 0.01$ between the TFE thickness and the local film thickness, independent of the effective gravity, viscosity, or thinning Regime. This value is consistent with previous measurements under standard gravity [nierstrasz1998marginal, yi_quantitative_2025] and slightly higher than the theoretical maximum of approximately 0.7 predicted by current models of marginal regeneration [mysels1959soap, gros2021marginal, cantat_drainage_2025]. The dynamics of these thin film elements depend on the effective gravity. In Regime 1, at low effective gravity, the film exhibits a strong thickness gradient, and TFEs stop at the boundary where their thickness equals the local film thickness. The central region is progressively eroded, producing a gradual reduction of the thickness gradient with a factor of 0.87 from the film center to the edge. In Regime 2, where TFEs populate the film from the start, the same proportionality factor of 0.87 is observed directly between the center and the boundary. In both regimes, these observations are modulated by inertial effects when $h_{R}$ is initially sufficiently thick.
@@
Although effective gravity modifies the overall thinning patterns—leading to two distinct regimes and an inertia-to-viscous transition—it does not alter the fundamental expression for the linear liquid flux exchanged between the film and the meniscus. However, gravity affects the meniscus size, and consequently the magnitude of the flux. Increasing rotation speed or effective gravity reduces the meniscus radius, thereby enhancing the flux and accelerating film drainage. The scaling follows an exponent of $3/2$ with rotation speed (or equivalently $3/4$ with effective gravity), consistent with theoretical predictions. Geometric effects associated with the frame shape may alter this scaling exponent [vigna_flowing_2025]. A crossover could occur if the meniscus radius $r_{\rm m}$ becomes comparable to the film thickness hh—a situation expected only at very high rotation speeds ($\omega\approx 3600$ rad·s-1, or $\tilde{g}\approx 200$), well beyond the current experimental stability limits. Finally, note that the long-time scaling for marginal regeneration–driven thinning, $h_{R}(t)\propto t^{-2/3}$, decreases more rapidly than the $t^{-1/2}$ scaling expected for thinning by internal flow (Eq. (2)), explaining why marginal regeneration remains the dominant mechanism even at very long times.

### v2

@@ III.3 Thickness measurements
In Fig. 4(a), we present typical time evolutions of full thickness profiles $h(r,t)$ in Regime 1, illustrating the coexistence of the two regions until the outer region fully invades the center.
[…]
In Fig. 4(b), we plot $h_{\rm R}$, $h_{\rm 0}$ and $h_{\rm b}$ as functions of time for the same dataset. Since the thickness profile in the central region remains steady, $h_{\rm 0}$ stays constant up to $t^{\star}$, the time at which the central zone disappears. We observe that both $h_{\rm R}$ and $h_{\rm b}$ decrease over time. Moreover, a proportionality relation is verified between these two quantities: as shown in Fig. 4(c), plotting $h_{\rm b}$ against $h_{\rm R}$ reveals a proportionality with a coefficient of approximately 0.87, which holds even at early times.
In Fig. 5(a), we plot these two quantities for the solution with viscosity $\eta=1$ mPa·s and for several rotation speeds, thus covering both Regimes 1 and 2. The proportionality relation between $h_{\rm b}$ and $h_{\rm R}$ holds at the lowest rotation speeds, with a proportionality coefficient of $0.87\pm 0.01$, obtained from a best fit over all corresponding data. At higher rotation speeds, a deviation from this proportionality appears at early times, when the film is relatively thick. We verified that these behaviors are consistent across all tested viscosities. We define $h_{\rm c}$ as the critical thickness below which the proportionality is recovered, and we find that $h_{\rm c}$ decreases with increasing rotation speed. In Fig. 5(b), we present an occurrence diagram distinguishing experiments that exhibit early-time deviations from proportionality. A comparison with the Regime diagram in Fig. 2(c) reveals a strong correlation: full proportionality is observed primarily in Regime 1, while partial proportionality systematically occurs in Regime 2, as well as in a few high-viscosity cases at the highest rotation speeds within Regime 1.
@@
First, when proportionality is observed, the coefficient of 0.87 is consistent with the value reported by Nierstrasz et al. [42] for vertical films under standard gravity. This supports the interpretation that TFEs nucleate along the frame edge, detach, and migrate radially inward under effective buoyancy, ultimately settling where the film thickness matches that of the TFEs (see the nucleation along the frame edge in Fig. 1(c), as well as the inward migration indicated by the arrows in panels (b) and (c) of the same figure). This scenario assumes that the TFE migration time is short compared to the overall thinning time, which is always the case in our observations. Our findings validate this mechanism even under centrifugal acceleration up to almost 100 times standard gravity. The gravity-independence of the 0.87 ratio reinforces the scenario proposed by Gros et al. [23], which attributes the selection mechanism to capillary suction and Marangoni stresses, rather than gravitational forces.
@@
which is the equation of an oscillator around the position where $h=h_{\rm TFE}$ as expressed in the third term. The first term corresponds to the inertial force, with a patch mass of $\rho h_{\rm TFE}\pi\ell_{\rm{TFE}}^{2}/4$ and added mass factor $\alpha$. The second term is a damping force with contributions from surface viscosity $\eta_{\rm s}$, bulk viscosity $\eta$, and air friction (with viscosity $\eta_{\rm{air}}$), such that the total two-dimensional viscosity is $\eta^{2D}=2\eta_{\rm s}+\eta h+k_{\rm{air}}\eta_{\rm{air}}\ell_{\rm{TFE}}$, where $k_{\rm{air}}$ is a geometric prefactor [35]. When damping dominates, TFEs directly reach their equilibrium position and $h_{\rm TFE}=h_{\rm b}$. This is consistent with observations in Regime 1 and Regime 2 when $h_{\rm R}<h_{\rm c}$. In contrast, during the early stages of Regime 2 (when $h_{\rm R}>h_{\rm c}$), inertia dominates, and TFEs overshoot their equilibrium and reach the film center.
@@
The diameter of TFE $\ell_{\rm{TFE}}$ has been measured in all experiments, both at the boundary between the two regions in Regime 1 and at the film center in Regime 2 (when $h_{\rm R}$ reaches $h_{\rm c}$). Figure 6(b) shows that $\ell_{\rm{TFE}}$ appears to be independent of $\omega$, and weakly dependent of $\eta$ only. In Eq. (7), we therefore consider $\ell_{\rm{TFE}}$ as a constant, approximately equal to $2$ mm. A detailed space–time analysis of $\ell_{\rm TFE}$ is beyond the scope of the present study.
Figure 6: (a) $h_{\rm R}(t=0)$, $h_{\rm c}$ (when a deviation is observed), and $h_{\rm R}(t=t_{\rm f})$ as functions of $\omega$, where $t_{\rm f}$ denotes the final measurement time. Data from all experiments and viscosities are shown with opacity encoding. (b) $\ell_{\rm TFE}$ as a function of $\omega$, showing that this quantity remains approximately constant. Measurements are taken when $h_{\rm R}=h_{\rm c}$, while TFEs are located at the film center. Data include all experiments in which a deviation from proportionality is observed.
@@ III.4 Thinning dynamics
As illustrated by the example in Fig. 4(b), internal flow described by Eq. (2), in blue dashed line, is too slow to account for the experimentally observed thinning dynamics, indicating that marginal regeneration is the dominant thinning mechanism in this case. This conclusion remains valid even at the highest effective gravities, as detailed in the Discussion and supported by the analysis that follows.
Assuming thinning driven by marginal regeneration, the expression for the linear flux $q$, as given in Eq. (1)—previously validated under standard gravity [32, 54, 11]—suggests a correlation between the rate of thinning $\textrm{d}h_{\rm R}/\textrm{d}t$ and the film thickness $h_{\rm R}$. Approximating the film volume as $\pi R^{2}h_{\rm R}$ and the liquid loss rate as $2\pi Rq$, we expect, following [32, 54, 11]:
@@
In Fig. 7(b), we explore the influence of rotation speed by plotting $\lvert\mathrm{d}h_{\rm R}/\mathrm{d}t\rvert$ versus $h_{\rm R}$ at fixed viscosity. Each dataset again follows a power law with an exponent close to $5/2$. Deviations appear when $h_{\rm R}>h_{\rm c}$, as shown by the triangles markers in Fig. 7(c) where the proportionality between $h_{\rm R}$ and $h_{\rm b}$ no longer holds; these data are excluded from the analysis that follows. As expected, higher rotation speeds lead to faster thinning. A power-law fit yields $\lvert\mathrm{d}h_{\rm R}/\mathrm{d}t\rvert\propto\omega^{1.31\pm 0.33}$. Although $\omega$ does not appear explicitly in Eq. (8), it influences the meniscus radius $r_{\rm m}$, which we assume to scale with the effective capillary length, $r_{\rm m}\propto\sqrt{\gamma/(\rho R\omega^{2})}$. Substituting this relation in Eq. (8) gives the thinning law:
$\frac{\textrm{d}h_{\rm R}}{\textrm{d}t}\propto-\frac{\omega^{3/2}\rho^{3/4}\gamma^{1/4}}{\eta\,R^{1/4}}\,h_{R}^{5/2}.$ (9)
Noting $K\propto{\omega^{3/2}\rho^{3/4}\gamma^{1/4}}/({\eta\,R^{1/4}})$ the proportionality constant between $\lvert\textrm{d}h_{R}/\textrm{d}t\rvert$ and $h_{R}^{5/2}$, the predicted dependence $K\propto\omega^{3/2}$ agrees well with the empirical exponent $1.31\pm 0.33$. When $\omega^{-3/2}\,\lvert\mathrm{d}h_{\rm R}/\mathrm{d}t\rvert$ is plotted against $h_{\rm R}$ for the data in Fig. 7(b), all curves collapse onto a single master curve (see inset), confirming the predicted $\omega$ scaling.
Overall, the expression for $K$ is validated over 3 orders of magnitude in Fig. 7(c), where the effects of both viscosity and rotation speed were independently tested. These results confirm thinning driven by marginal regeneration and that Eq. (1) remains valid even under high-gravity conditions. By integrating Eq. (9) over time, we obtain a formula analogous to Eq. (2) for internal flow, but describing thinning driven by marginal regeneration:
$h_{\rm R}(t)=\frac{h_{i}}{\left(1+\frac{3}{2}Kh_{i}^{3/2}(t-t_{i})\right)^{2/3}},$ (10)
where $h_{i}$ is the initial thickness at time $t_{i}$. Testing this expression on the data from Fig. 4(b) using the corresponding value of $K$ from Fig. 7(c) shows good agreement. Furthermore, Eq. (10) exhibits the asymptotic scaling $h_{\rm R}\propto t^{-2/3}$, consistent with previous studies under standard gravity [32, 37].
@@ III.5 Discussion
At early stages, in the absence of marginal regeneration, centrifugal acceleration likely induces an initial stretching of the film rather than thinning through internal flow, although existing interfacial elasticity models fail to reproduce the observed profiles. This stretching occurs only in the central region and at low effective gravity, in Regime 1, before erosion by marginal regeneration begins. In Regime 2, TFEs populate the entire film as soon as imaging begins, suggesting that either the initial erosion of the central zone is too fast to be captured, or that the film is sufficiently flat for TFEs to migrate directly toward the center. Interestingly, films remain very stable, particularly in the absence of glycerol. Assuming an average film thickness of a few micrometers, the surface tension difference between the center and the frame edge can be estimated from Eq. (3) as $\Delta\gamma=\rho hR^{2}\omega^{2}/4\sim 7$ mN·m-1, smaller than but comparable to the surface tension of the surfactant solution. This difference may explain why the system approaches the limit of film stability reported in previous studies [17].
The presence or absence of a central zone remains one of the open questions of this study. The transition between Regimes 1 and 2 occurs at effective gravities of approximately 3–5 times standard gravity, with only a weak dependence on viscosity. Investigating the transient stage during which the film is brought into solid-body rotation is therefore crucial for understanding the early-stage dynamics and, in particular, the fate of the central zone, as well as how marginal regeneration operates during this period—potentially coupled with interface elasticity and, to a lesser extent, internal flow. Future work could address this issue by controlling the duration of this transient stage—about 1 s in the present study—for instance by progressively ramping up the rotation speed.
[…]
@@
Across all Regimes and viscosities, thinning is consistently associated with the nucleation of TFEs in the vicinity of the meniscus. These elements are thinner than the surrounding film, resulting in a net mass loss as they migrate inward. On average, we find a proportionality factor of $0.87\pm 0.01$ between the TFE thickness and the local film thickness, independent of the effective gravity, viscosity, or thinning Regime. This value is consistent with previous measurements under standard gravity [42, 58] and slightly higher than the theoretical maximum of approximately 0.7 predicted by current models of marginal regeneration [41, 23, 11]. The dynamics of these thin film elements depend on the effective gravity. In Regime 1, at low effective gravity, the film exhibits a strong thickness gradient, and TFEs stop at the boundary where their thickness equals the local film thickness. The central region is progressively eroded, producing a gradual reduction of the thickness gradient with a factor of 0.87 from the film center to the edge. In Regime 2, where TFEs populate the film from the start, the same proportionality factor of 0.87 is observed directly between the center and the boundary. In both regimes, these observations are modulated by inertial effects when $h_{R}$ is initially sufficiently thick.
@@
Although effective gravity modifies the overall thinning patterns—leading to two distinct regimes and an inertia-to-viscous transition—it does not alter the fundamental expression for the linear liquid flux exchanged between the film and the meniscus. However, gravity affects the meniscus size, and consequently the magnitude of the flux. Increasing rotation speed or effective gravity reduces the meniscus radius, thereby enhancing the flux and accelerating film drainage. The scaling follows an exponent of $3/2$ with rotation speed (or equivalently $3/4$ with effective gravity), consistent with theoretical predictions. Geometric effects associated with the frame shape may alter this scaling exponent [54]. A crossover could occur if the meniscus radius $r_{\rm m}$ becomes comparable to the film thickness hh—a situation expected only at very high rotation speeds ($\omega\approx 3600$ rad·s-1, or $\tilde{g}\approx 200$), well beyond the current experimental stability limits.